OpenStudy (anonymous):

A plane flew 1740 km in 6 hours with a tail wind. On the return trip, going into the wind, the plane flew 570 km in 3 hours. What is the plane's speed if there was no wind? What is the speed of the wind? Write the system of equations that would be used to solve this problem.

OpenStudy (anonymous):

1740km is 6 hours So 870 in 3 hours Then 570 in 3 hours on the way back The difference is 300km per 3 hours but because it's acting in opposite directions it works out as 150 kn per 3 hours so the wind speed is 50km/ hour The planes speed without wind would be the average so 720 km per 3 hours so 240km/hour

OpenStudy (anonymous):

Wind's speed (y): 6(x + y) = 1,740 x + y = 290 y = 290 - x 3(x - y) = 570 x - y = 190 y = x - 190 Plane's speed (x): 290 - x = x - 190 2x = 480 x = 240 Wind's speed: = 290 - 240 = 50

OpenStudy (anonymous):

theyre both right